he comparative balance sheet of Olson-Jones Industries Inc. for December 31, 20Y2 and 20Y1, is as follows: 1 Dec. 31, 20Y2 Dec. 31, 20Y1 2 Assets 3 Cash $488.00 $163.00 4 Accounts receivable (net) 283.00 203.00 5 Inventories 171.00 110.00 6 Land 396.00 455.00 7 Equipment 225.00 171.00 8 Accumulated depreciation-equipment (56.00) (26.00) 9 Total assets $1,507.00 $1,076.00 10 Liabilities and Stockholders’ Equity 11 Accounts payable (merchandise creditors) $173.00 $164.00 12 Dividends payable 26.00 13 Common stock, $1 par 98.00 47.00 14 Paid-in capital: Excess of issue price over par—common stock 254.00 122.00 15 Retained earnings 956.00 743.00 16 Total liabilities and stockholders’ equity $1,507.00 $1,076.00 The following additional information is taken from the records: A. Land was sold for $122. B. Equipment was acquired for cash. C. There were no disposals of equipment during the year. D. The common stock was issued for cash. E. There was a $313 credit to Retained Earnings for net income. F. There was a $100 debit to Retained Earnings for cash dividends declared. A. Prepare a statement of cash flows, using the indirect method of presenting cash flows from operating activities. Refer to the Labels and Amount Descriptions list provided for the exact wording of the answer choices for text entries. Be sure to complete the heading of the statement. In the operating activities section, use the minus sign to indicate cash outflows, decreases in cash and a net cash out flow, if required. In the investing and financing activities section, use a minus sign only to indicate a NET cash outflow for the section. B. Was Olson-Jones’s net cash flow from operations more or less than net income? What is the source of this difference?
Answer #1

A.

Olson-Jones Industries Inc Statement of Cash Flows (by Indirect Method) For the Year Ended December 31, 20Y2 Cash flows from Operating activities: Net Income Adjustments to reconcile net income to net cash from operating activities $313 Add: (Gain)/Loss on Sale of Land Add: Depreciation Less: Increase in Inventories Less: Increase in Accounts Receivable Add: Increase in Accounts Payable -$63 $30 -$61 $80 $9 $148 Net Cash from Operating Activities [A] Cash flows from Investing activities: Sale of Land Purchase of Equipment Issue of Common Stock Issue of Common Stock-Price over Payment of Dividends Net Cash used in Investing Activities [B] $122 $54 $51 $132 $74 $177 Par Cash flows from Financing activities: Net Cash from Financing Activities [C] $0 Net Increase in Cash during the year [A+B+C] Add: Cash, January 1, 20Y2 Add: Cash, December 31, 20Y2 $325 $163 $488Workings: Gain/Loss on Sale of Land (Gain)/Loss on Sale of Land = Cost of Land-Sale Price of Land - (455-396) 122 S (63 Depreciation Depreciation = Current Year Balance-Previous Year Balance = 56-26 30 Increase in Inventories Previous Year Balance-Current Year Balance = 110-171 = s (61) Increase in Inventories Increase in Accounts Receivable Increase in Accounts Receivable Previous Year Balance-Current Year Balance - 203-283 $ (80) Increase in Accounts Payable Increase in Accounts Payable - Current Year Balance Previous Year Balance - 173-164 9Purchase of Equipment Purchase of Equipment Current Year Balance-Previous Year Balance - 225 171 - 54 Issue of Common Stock Issue of Common Stock = Current Year Balance-Previous Year Balance = 98-47 $51 Issue of Common Stock-Price over Par Issue of Common Stock-Price over Par -Current Year Balance-Previous Year Balance = 254-122 = $132 Payment of Dividends Payment of Dividends Dividend Declared - Dividend Payable 100-26 $ 74

B. Net Cash flow from Operating Activities is lower than Net Income by $165. The reason for the same is the changes in working capital, profit/loss on sale of asset and depreciation. The first one is not considered in Income Statement, but for preparation of Cash Flow as it is a cash item, we have to consider the difference between the opening and closing amounts. The latter two being non-cash items which were included in Income Statement are excluded from Cash Flow Statement.